A certain trip took 40 min to complete. If half the distance was 100 and the other half at 60, what's the distance traveled during the trip?
a) 68 b) 50

Answer by checkley77(12569)

D=RT
T=D/R
2/3=(D/2)/100+(D/2)/60
2/3=D/2*1/100+D/2*1/60
2/3=D/200+D/120
2/3=(120D+200D)/(200*120)
2/3=320D/24,000
3*320D=2*24,000
960D=48,000
D=48,000/960
D=50 MILES IS THE DISTANCE OF THE TRIP.